Double your wealth. Kant Miss Company is promising its investors that it will double their money every 4 years. What annual rate is Kant Miss​ promising? Is this investment a good​ deal? If you invest ​$400 now and Kant Miss is able to deliver on its​ promise, how long will it take your investment to reach ​$28,000​?
Answer #1


solution: a) FV=PV*(1+i)^n 2=1*(1+i)^4 (1+i)^4=2/1 (1+i)^4=2 1+i=2^(1/4) 1+1=1.189207 i=1.189207-1 FV Future value PV Presentb) FV=PV*(1+r)^n 28000=400*(1.1892071)^n (1.1892071)^n=28000/400 (1.1892071)^n=70 n log 1.1892071=log 70 n=log 70/log 1.18920